OpenStudy (anonymous):

The slope of TV is ½ and V (5,1) . Find the coordinate of point T that would hold this information true.

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

can't wait to see this there are an infinite number of possible answers as the question is stated is there something missing?

OpenStudy (anonymous):

There's 4 answer choices.

OpenStudy (anonymous):

name one and we can check it

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

(10,3)

OpenStudy (anonymous):

the slope of the line containing \((5,1)\) and \((10,3)\) is \[\frac{3-1}{10-5}=\frac{2}{5}\] which is not \(\frac{1}{2}\) so that is not the answer

OpenStudy (anonymous):

Ok, thanks:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Should I name another?

OpenStudy (anonymous):

presumably one pair will make the slope \(\frac{1}{2}\) when it is, that is the correct one it is usually C or sometimes D

OpenStudy (anonymous):

Hey, Ari

OpenStudy (aripotta):

oh hai

OpenStudy (anonymous):

How are you doing?

OpenStudy (anonymous):

Oh, satellite is here….gotta do math

OpenStudy (aripotta):

i'm ok. but the day just started and i can already tell it's gonna be bad.

OpenStudy (anonymous):

How come?

OpenStudy (aripotta):

my family is working on getting the house together before christmas. and whenever we do that, grandma starts yelling and really nothing gets done. but anyway, do you still need help on this?

OpenStudy (anonymous):

No, don';t think so. Thanks though,

OpenStudy (anonymous):

I'll figure it out.

OpenStudy (anonymous):

I'll just go by the example that the mod gave me.

OpenStudy (aripotta):

alright, well satellite's there for ya :P

OpenStudy (anonymous):

True. Thanks though!!!!

OpenStudy (anonymous):

I got (-1,-2)